OpenStudy (anonymous):

Help with Algebra yo? Which of the following functions represents the table above? A.y = 3x3 B.y = 3x C.y =√ (x) D.y = 3x2

OpenStudy (anonymous):

This is what the graph looks like: X 0 1 2 3 4 Y 0 3 24 81 192

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Have you tried plugging in the value for x in the table for each of the answers? For example, for x=1 in the table. Plugging into each of answers we are looking for y=3 Answer A: y= 3(x)^3 = 3(1)=3 This is a possible correct answer Answer B: y=3x = 3(1)= 3 This is a possible correct answer. Answer C: y=sqrt(x) = sqrt(1)=1 This the wrong. So this answer is no good. Answer D: y=3(x)^2 = 3(1)^2 = 3(1)=3 This is a possible correct answer. So answers A, B and D are possible correct answers. So try another value for x. For example x=2 in the table. We are looking for y=24 Answer A: y=3(x)^3 = I'll let you figure this one. Answer B; y= 3(x)= 3(2) =6 This is wrong. So this answer is no good. Answer C: We already said this answer was no good. Answer D: y= 3(x)^2 = 3(2)^2= 3(4)=12 This is wrong so this answer is no good.